# Huppert’s conjecture for finite simple exceptional groups of Lie type

### v1

@@ 5.3. Reduction to finite simple groups of Lie type in characteristic $p$
$(a)$ Case $S\cong{\mathrm{L}}_{n}^{\epsilon}(p^{b}),$ with $n\geq 2.$ We have $bn(n-1)=48f.$ If $n=2$ then $b=24f$ so $S={\mathrm{L}}_{2}(q^{24})$ and hence $S$ has a character of degree $q^{24}+1.$ Obviously this degree divides no degree of $G$ since $q^{24}+1\nmid|{\rm F}_{4}(q)|.$ Next if $n=3$ then $b=8f$ and so $S={\mathrm{L}}_{3}^{\epsilon}(q^{8}).$ By [4, $(13.8)$], $S$ possesses a unipotent character parametrised by the partition $(1,2)$ of degree $q^{8}(q^{8}+\epsilon 1).$ As $q^{8}+1$ does not divide the order of $H$ so the case when $\epsilon=+$ cannot happen. Now assume $\epsilon=-.$ Then $q^{8}(q^{8}+\epsilon 1)=q^{8}(q^{8}-1)=q^{8}\Phi_{1}\Phi_{2}\Phi_{4}\Phi_{8},$ which is coprime to $\ell_{3}\ell_{4},$ so $q^{8}(q^{8}-1)$ must be one of the degrees given in case $(iv)$ of Lemma 3.1. However by comparing the power of $q,$ we see that this case is impossible. If $n=4,$ then $b=4f$ so $S={\mathrm{L}}_{4}^{\epsilon}(q^{4}).$ In this case, the unipotent character parametrized by the partition $(2,2)$ has degree $q^{8}(q^{8}+1).$ As above, this degree does not belong to ${\mathrm{cd}}(G).$ Thus we can assume that $n\geq 5.$ By Table 2, $S$ possesses a unipotent character $\chi\neq{\mathrm{St}}_{S}$ with $\chi(1)_{p}=p^{b(n-1)(n-2)/2}.$ By Lemma 3.1$(vi)$, we have $b(n-1)(n-2)/2\leq 16f$ if $p$ is odd and $b(n-1)(n-2)/2\leq 13f$ if $p$ is even. Assume first that $p$ is even. Multiplying both sides of the inequality $b(n-1)(n-2)/2\leq 13f$ by $n,$ we obtain $bn(n-1)(n-2)/2=24f(n-2)\leq 13nf,$ and so $24(n-2)\leq 13n.$ Thus $11n\leq 48$ so $n\leq 4,$ which is a contradiction. Thus $p$ is odd. Using the same argument with $b(n-1)(n-2)/2\leq 16f,$ we obtain that $n\leq 6.$ Hence $5\leq n\leq 6$ and $p$ is odd. Assume that $n=5.$ Then $5b=12f.$ We have that the primitive prime divisor $\ell_{4b}(p)$ exists and so as $\ell_{4b}(p)\mid|S|,$ we deduce that $\ell_{4b}(p)\mid|H|.$ Since $4b=48f/5>8f,$ we must have that $4b\mid 12f=5b,$ which is impossible. Now assume $n=6.$ Then $5b=8f$ and the primitive prime divisor $\ell_{6b}(p)\in\pi(S)\subseteq\pi(H)$ exists. As $6b>5b=8f$ we have that $6b\mid 12f,$ and hence $b\mid 2f,$ which implies that $5b=8f\mid 10f,$ which is impossible.
@@
If $n=4$ then $2b=3f$ so $b\geq 3.$ Hence $\ell_{6b}(p)\in\pi(S)\subseteq\pi(H),$ and $6b=9f>8f,$ we deduce that $6b=9f\mid 12f,$ which is impossible.
If $n=5$ then $25b=24f.$ We have that $10b=240f/25=48f/5>9f$ so $\ell_{10b}(p)\in\pi(S)\subseteq\pi(H),$ hence $10b\mid 12f,$ which implies that $5b\mid 6f$ and so $25b=24f\mid 30f,$ which is impossible.
Hence we can assume that $n\geq 6.$ By Table 2, $S$ possesses a nontrivial irreducible character $\chi\neq{\mathrm{St}}_{S}$ with $\chi(1)_{p}=p^{b(n-1)^{2}}/(2,p).$ Assume first that $p=2.$ Then $\chi(1)_{p}=p^{b(n-1)^{2}-1}$ and hence by 3.1$(vi),$ we obtain $bn(n-2)\leq b(n-1)^{2}-1\leq 13f.$ Multiplying both sides by $n$ and simplifying, we have $11n\leq 48$ so $n<5,$ which is a contradiction. Now assume $p>2.$ Then $\chi(1)_{p}=p^{b(n-1)^{2}}$ so by Lemma 3.1$(vi),$ we have that $b(n-1)^{2}\leq 16f.$ Multiplying both sides by $3,$ we obtain $3b(n-1)^{2}\leq 48f=2bn^{2}.$ After simplifying, we have $n(n-6)+3\leq 0,$ which is impossible as $n\geq 6.$
$(c)$ Case $S\cong{\mathrm{O}}_{2n}^{\epsilon}(q_{1}),$ where $q_{1}=p^{b},$ and $n\geq 4.$ We have $bn(n-1)=24.$ If $n=4,$ then $b=2f$ hence $q_{1}=q^{2}.$ Now if $S={\mathrm{O}}_{8}^{+}(q^{2}),$ then $S$ has a unipotent character $\chi$ labeled by the symbol in Table 2 with $\chi(1)=q^{14}\Phi_{8}^{2}$ and if $S={\mathrm{O}}_{8}^{-}(q^{2}),$ then $S$ has a unipotent character $\chi$ labeled by the symbol $\binom{1~{}3}{~{}-}$ with degree $\chi(1)={(q^{8}+1)(q^{6}-q^{2})}/{(q^{4}-1)}.$ However these degrees do not divide the order of $H.$ If $n=5,$ then $5b=6f.$ We have that $8b=48f/5>9f$ and so $\ell_{8b}(p)\in\pi(S)\subseteq\pi(H),$ so $8b\mid 12f=10b,$ which is impossible. If $n=6,$ then $5b=4f.$ We have that $8b=32f/5>6f$ and so $\ell_{8b}(p)\in\pi(S)\subseteq\pi(H),$ so $8b\mid 8f=10b$ or $8b\mid 12f=15b.$ However both cases are impossible. Thus we can assume that $n\geq 7.$ By Table 2, $S$ possesses a unipotent character $\chi\neq{\mathrm{St}}_{S}$ with $\chi(1)_{p}\geq p^{b(n-1)(n-2)}.$ By Lemma 3.1$(vi),$ we have $b(n-1)(n-2)\leq 16f.$ Multiplying both sides by $n,$ and simplifying, we obtain $8n\leq 48$ and hence $n\leq 6,$ which is a contradiction.
@@
$(e)$ Case $S\cong{\rm G}_{2}(q_{1}),$ where $q_{1}=p^{b}$. We have $6b=24f$ and so $b=4f.$ Thus $S={\rm G}_{2}(q^{4}),$ so $S$ has an irreducible character of degree $q^{24}-1$ by [5]. However this degree divides no degrees of $H.$
@@
$(g)$ For the remaining cases, we can argue as follows. We have that $|S|_{p}=p^{24f}.$ Lemma 3.1$(vi)$ yields that $\chi(1)_{p}\leq p^{16f},$ where $\chi$ is the unipotent character listed in Table 2. Using these two properties, we will obtain a contradiction. For instance, assume that $S\cong{\rm E}_{8}(p^{b}),$ where $b\geq 1.$ Then $120b=24f$ and hence $5b=f.$ By Table 2, $S$ possesses a unipotent character $\chi\neq{\mathrm{St}}_{S}$ with $\chi(1)_{p}\geq p^{91b}.$ By Lemma 3.1$(vi),$ we have $91b\leq 16f=80b,$ a contradiction.
@@
$(a)$ Case $S\cong{\mathrm{L}}_{n}^{\epsilon}(p^{b}),$ with $n\geq 2.$ We have $bn(n-1)=72f.$ If $n=2$ then $b=36f$ so $S={\mathrm{L}}_{2}(q^{36})$ and hence $S$ has a character of degree $q^{36}+1.$ However this is impossible as this number does not divide the order of $H.$ Hence we can assume that $n\geq 3.$ If $(n,q)=(6,2)$ or $(3,2),$ then obviously the equation $bn(n-1)=72f$ cannot happen. Thus $\ell_{\epsilon bn}(p)\in\pi(S)\subseteq\pi(H)$ always exists by Lemma 2.5, and so $bn\leq 18f.$ Multiplying both sides by $n-1$ and simplifying, we obtain that $n\geq 5.$ By Table 2, $S$ possesses a unipotent character $\chi\neq{\mathrm{St}}_{S}$ with $\chi(1)_{p}=p^{b(n-1)(n-2)/2}.$ By Lemma 3.2(vi) we have $b(n-1)(n-2)/2\leq 25f.$ Multiplying both sides of the latter inequality by $n$ and simplifying, we obtain that $n\leq 6.$ If $n=5,$ then $5b=18f$ and hence $4b=72f/5>14f$ so $\ell_{4p}(p)\in\pi(S)\subseteq\pi(H)$ exists and thus $4b\mid 18f=5b,$ which is impossible. Similarly if $n=6,$ then $5b=12f$ and hence $6b=72f/5>14f$ so $\ell_{6p}(p)\in\pi(S)\subseteq\pi(H)$ exists and thus $6b\mid 18f.$ Hence $12b\mid 36f=15b,$ a contradiction.
@@
$(b)$ Case $S\cong{\mathrm{S}}_{2n}(q_{1}),$ or ${\mathrm{O}}_{2n+1}(q_{1}),$ with $q_{1}=p^{b},$ $n\geq 2$ and $S\neq{\mathrm{S}}_{4}(2).$ We have $bn^{2}=36f.$ If $p=2$ and $2nb=6,$ then $n=3,b=1$ and so $bn^{2}=9=36f,$ a contradiction. Thus $\ell_{2bn}(p)\in\pi(S)\subseteq\pi(H)$ exists. Hence $2bn\leq 18f.$ Multiplying both sides by $n$ and simplifying, we obtain $n\geq 4.$ By Table 2, $S$ possesses a nontrivial irreducible character $\chi\neq{\mathrm{St}}_{S}$ with $\chi(1)_{p}\geq p^{bn(n-2)},$ and hence by Lemma 3.2$(vi)$, $bn(n-2)\leq 25f.$ Multiplying both sides by $n$ and simplifying, we have $11n\leq 72$ so $n\leq 6.$ If $n=4,$ then $4b=9f$ and hence $6b=54f/4>12f$ so $\ell_{6p}(p)\in\pi(S)\subseteq\pi(H)$ exists and thus $6b\mid 18f=8b,$ which is impossible. Similarly if $n=5,$ then $25b=36f$ and hence $10b=360f/25>14f$ so $\ell_{10p}(p)\in\pi(S)\subseteq\pi(H)$ exists and thus $10b\mid 18f.$ It follows that $20b\mid 36f=25b,$ which is impossible. Thus $n=6$ and so $b=f.$ By [4], $S$ possesses a unipotent character $\chi^{\alpha}$ labeled by the symbol $\alpha=\binom{0~{}1~{}6}{~{}-~{}}$ with degree $\chi^{\alpha}(1)=(q^{6}-1)(q^{6}-q)/(2(q+1)).$ Obviously $\chi^{\alpha}(1)$ is nontrivial and different from $|H|_{p}$ and furthermore it is coprime to $\ell_{i}$ for $i=1,2,3,$ which contradicts Lemma 3.2$(i)$.
@@
$(c)$ Case $S\cong{\mathrm{O}}_{2n}^{\epsilon}(q_{1}),$ where $q_{1}=p^{b},$ and $n\geq 4.$ We have $bn(n-1)=36f.$ By Table 2, $S$ possesses a unipotent character $\chi\neq{\mathrm{St}}_{S}$ with $\chi(1)_{p}\geq p^{b(n-1)(n-2)}.$ By Lemma 3.2$(vi)$, we have $b(n-1)(n-2)\leq 25f.$ Multiplying both sides by $n,$ and simplifying, we obtain $n\leq 6.$ If $n=4,$ then $b=3f$ and hence $q_{1}=q^{3}.$ By [4], $S$ possesses a unipotent character $\varphi$ labeled by the symbol $\binom{1~{}n}{0~{}1}$ when $\epsilon=+$ and $\binom{0~{}1~{}n}{~{}~{}1~{}~{}}$ when $\epsilon=-$ of degree $(q_{1}^{2n}-q_{1}^{2})(q_{1}^{2}-1)=(q_{1}^{8}-q_{1}^{2})/(q_{1}^{2}-1)=q^{6}\Phi_{9}\Phi_{18}.$ It follows that $\varphi(1)\in{\mathrm{cd}}(L)-\{1,q^{36}\}$ and $(\varphi(1),\ell_{2}\ell_{3})=1.$ By Lemma 3.2$(ii)$, $\varphi(1)$ must be one of the degrees in $(ii).$ However both cases are impossible by comparing the power on $q.$ If $n=5,$ then $5b=9f$ and hence $8b=72f/5>14f$ so $\ell_{8p}(p)\in\pi(S)\subseteq\pi(H)$ exists and thus $8b\mid 18f=10b,$ which is impossible. Thus $n=6$ and so $5b=6f.$ As in case $n=4,$ $S$ possesses a unipotent character $\varphi$ of degree $(q_{1}^{12}-q_{1}^{2})/(q_{1}^{2}-1)=q_{1}^{2}(q_{1}^{10}-1)/(q_{1}^{2}-1)=q_{1}^{2}(q^{12}-1)/(q_{1}^{2}-1).$ It follows that $\varphi(1)\in{\mathrm{cd}}(L)-\{1,q^{36}\}$ and $(\varphi(1),\ell_{1}\ell_{3})=1,$ contradicting Lemma 3.2$(iii)$.
@@
$(d)$ Case $S\cong{\rm G}_{2}(p^{b}),$ where $b\geq 1.$ We have $6b=36f.$ By [5], $S$ possesses an irreducible character of degree $p^{6b}-1=q^{36}-1.$ However $q^{36}-1$ does not divide $|H|,$ a contradiction.
@@
$(e)$ Case $S\in\{{}^{2}{\rm B}_{2}(q_{1}^{2}),{}^{2}{\rm G}_{2}(q_{1}^{2}),$ where $q_{1}^{2}=2^{2n+1},3^{2n+1},$ respectively with $n\geq 1.$ In these cases, the equation $|S|_{p}=|H|_{p}$ cannot happen.
@@
$(f)$ Case $S\cong{}^{2}{\rm F}_{4}(q_{1}^{2}),$ where $q_{1}^{2}=2^{2n+1},$ with $n\geq 1.$ Then $12(2n+1)=36f.$ We have that $\ell_{12(2n+1)}(2)\in\pi(S)\subseteq\pi(H),$ which implies that $12(2n+1)=36f\leq 18f,$ which is impossible.
@@
$(i)$ Case $S\cong{}^{3}{\rm D}_{4}(p^{b})$. Then $12b=36f.$ We have $\ell_{12b}(p)\in\pi(S)$ but $\ell_{12b}(p)=\ell_{36f}(p)\not\in\pi(H).$
@@
$(k)$ Case $S\cong{\rm E}_{7}(p^{b})$. Then $63b=36f$ and hence $7b=4f.$ By Table 1, $S$ possesses a unipotent character $\varphi\neq{\mathrm{St}}_{S}$ with $\varphi(1)_{p}=p^{46b}.$ By Lemma 3.2$(vi)$, we obtain that $46b\leq 25f,$ which implies that $184b=4\cdot 46b\leq 100f=175b,$ which is impossible.
$(l)$ Case $S\cong{\rm E}_{8}(p^{b})$. Then $120b=36f$ and hence $10b=3f.$ By Table 1 and Lemma 3.2$(vi)$, we have that $91b\leq 25f.$ It follows that $273b\leq 25\cdot 3f=250b,$ a contradiction.
$(m)$ Case $S\cong{\rm E}_{6}(p^{b})$. Then $36b=36f$ and hence $b=f.$ By [4, 13.9], $S$ has a unipotent character $\varphi$ labeled by the symbol $\phi_{6,1}$ with degree $q\Phi_{8}\Phi_{9}.$ We have that $\varphi(1)\in{\mathrm{cd}}(L)-\{1,|H|_{p}\}$ is coprime to $\ell_{1}\ell_{2}\ell_{4},$ which contradicts Lemma 3.2$(i)$.
The arguments for the remaining simple groups ${\rm E}_{6}(q),{\rm E}_{7}(q)$ and ${\rm E}_{8}(q)$ are similar, so we will skip the details.

### v3

@@ 5. Verifying Step 2: $G^{\prime}/M$ is isomorphic to $H$
$(a)$ Case $S\cong{\mathrm{L}}_{n}^{\epsilon}(p^{b}),$ with $n\geq 2.$ Here $bn(n-1)=48f.$ If $n=2$ then $b=24f$, so $S={\mathrm{L}}_{2}(q^{24})$. Hence $S$ has a character of degree $q^{24}+1.$ Obviously this degree divides no degree of $G$ since $q^{24}+1\nmid|{\rm F}_{4}(q)|.$ If $n=3$ then $b=8f$ and so $S={\mathrm{L}}_{3}^{\epsilon}(q^{8}).$ By [4, $(13.8)$], $S$ possesses a unipotent character parametrized by the partition $(1,2)$ of degree $q^{8}(q^{8}+\epsilon 1).$ Since $q^{8}+1$ does not divide the order of $H$, the case when $\epsilon=+$ cannot occur. Now assume $\epsilon=-.$ Then $q^{8}(q^{8}+\epsilon 1)=q^{8}(q^{8}-1)=q^{8}\Phi_{1}\Phi_{2}\Phi_{4}\Phi_{8},$ which is coprime to $\ell_{3}\ell_{4},$ so $q^{8}(q^{8}-1)$ must be one of the degrees given in case $(iv)$ of Lemma 3.2. However by comparing the power of $q,$ we see that this case is impossible. If $n=4,$ then $b=4f$ so $S={\mathrm{L}}_{4}^{\epsilon}(q^{4}).$ In this case, the unipotent character parametrized by the partition $(2,2)$ has degree $q^{8}(q^{8}+1).$ As above, this degree does not belong to ${\mathrm{cd}}(G).$ Thus we assume that $n\geq 5.$ By Table 2, $S$ possesses a unipotent character $\chi\neq{\mathrm{St}}_{S}$ with $\chi(1)_{p}=p^{b(n-1)(n-2)/2}.$ By Lemma 3.2$(vi)$, if $p$ is odd, then $b(n-1)(n-2)/2\leq 16f$; else $b(n-1)(n-2)/2\leq 13f$. Assume first that $p$ is even. Multiplying both sides of the inequality $b(n-1)(n-2)/2\leq 13f$ by $n,$ we obtain
$bn(n-1)(n-2)/2=24f(n-2)\leq 13nf,$
and so $24(n-2)\leq 13n.$ Thus $11n\leq 48$ so $n\leq 4,$ a contradiction. Thus $p$ is odd. Using the same argument with $b(n-1)(n-2)/2\leq 16f,$ we obtain that $n\leq 6.$ Hence $5\leq n\leq 6$ and $p$ is odd. Assume that $n=5.$ Then $5b=12f.$ Since the primitive prime divisor $\ell_{4b}(p)$ exists, $\ell_{4b}(p)\mid|S|,$ and hence $\ell_{4b}(p)\mid|H|.$ Since $4b=48f/5>8f$, $4b\mid 12f=5b,$ which is impossible. Assume $n=6.$ Then $5b=8f$ and the primitive prime divisor $\ell_{6b}(p)\in\pi(S)\subseteq\pi(H)$ exists. As $6b>5b=8f$, $6b\mid 12f,$ and hence $b\mid 2f$. It follows that $5b=8f\mid 10f,$ which is impossible.
@@
${q_{1}^{3}(q_{1}^{6}-1)}/({q_{1}^{2}-1})={q^{8}(q^{16}-1)}/({q^{16/3}-1}).$
However $\chi(1)$ is not a degree of $G,$ since $\ell_{16}(q)\mid\chi(1)$ but $\ell_{16}(q)\nmid|H|.$
If $n=4$ then $2b=3f$ so $b\geq 3.$ Hence $\ell_{6b}(p)\in\pi(S)\subseteq\pi(H),$ and $6b=9f>8f,$ and we deduce that $6b=9f\mid 12f,$ which is impossible.
If $n=5$ then $25b=24f.$ Now $10b=240f/25=48f/5>9f$ so $\ell_{10b}(p)\in\pi(S)\subseteq\pi(H),$ hence $10b\mid 12f,$ which implies that $5b\mid 6f$ and so $25b=24f\mid 30f,$ which is impossible.
Assume $n\geq 6.$ By Table 2, $S$ possesses a nontrivial irreducible character $\chi\neq{\mathrm{St}}_{S}$ with $\chi(1)_{p}=p^{b(n-1)^{2}}/(2,p).$ Assume that $p=2.$ Then $\chi(1)_{p}=p^{b(n-1)^{2}-1}$. By Lemma 3.2$(vi),$
$bn(n-2)\leq b(n-1)^{2}-1\leq 13f.$
Multiplying both sides by $n$ and simplifying, $11n\leq 48$ so $n<5,$ a contradiction. Now assume $p>2.$ Then $\chi(1)_{p}=p^{b(n-1)^{2}}$. By Lemma 3.2$(vi),$ $b(n-1)^{2}\leq 16f.$ Multiplying both sides by $3,$
$3b(n-1)^{2}\leq 48f=2bn^{2}.$
After simplifying, $n(n-6)+3\leq 0,$ which is a contradiction as $n\geq 6.$
$(c)$ Case $S\cong{\mathrm{O}}_{2n}^{\epsilon}(q_{1}),$ where $q_{1}=p^{b},$ and $n\geq 4.$ Here $bn(n-1)=24.$ If $n=4,$ then $b=2f$ hence $q_{1}=q^{2}.$ If $S={\mathrm{O}}_{8}^{+}(q^{2}),$ then $S$ has a unipotent character $\chi$ labeled by the symbol in Table 2 with $\chi(1)=q^{14}\Phi_{8}^{2}$. If $S={\mathrm{O}}_{8}^{-}(q^{2}),$ then $S$ has a unipotent character $\chi$ labeled by the symbol $\binom{1~{}3}{~{}-}$ with degree $\chi(1)={(q^{8}+1)(q^{6}-q^{2})}/{(q^{4}-1)}.$ However, these degrees do not divide $|H|.$ If $n=5,$ then $5b=6f.$ Since $8b=48f/5>9f$, $\ell_{8b}(p)\in\pi(S)\subseteq\pi(H),$ so $8b\mid 12f=10b,$ which is impossible. If $n=6,$ then $5b=4f.$ Now $8b=32f/5>6f$ and so $\ell_{8b}(p)\in\pi(S)\subseteq\pi(H),$ thus $8b\mid 8f=10b$ or $8b\mid 12f=15b.$ Both cases are impossible. Thus we can assume that $n\geq 7.$ By Table 2, $S$ possesses a unipotent character $\chi\neq{\mathrm{St}}_{S}$ with $\chi(1)_{p}\geq p^{b(n-1)(n-2)}.$ By Lemma 3.2$(vi),$ $b(n-1)(n-2)\leq 16f.$ Multiplying both sides by $n,$ and simplifying, $8n\leq 48$ and hence $n\leq 6,$ a contradiction.
@@
$(e)$ Case $S\cong{\rm G}_{2}(q_{1}),$ where $q_{1}=p^{b}$. Here $6b=24f$ and so $b=4f.$ Thus $S={\rm G}_{2}(q^{4}),$ so $S$ has an irreducible character of degree $q^{24}-1$ by [5]. This degree divides no degrees of $H.$
@@
$(h)$ For the remaining cases, we argue as follows. In all cases $|S|_{p}=p^{24f}.$ Lemma 3.2$(vi)$ yields that $\chi(1)_{p}\leq p^{16f},$ where $\chi$ is the unipotent character listed in Table 2. Using these two properties, we obtain a contradiction. For instance, assume that $S\cong{\rm E}_{8}(p^{b}),$ where $b\geq 1.$ Then $120b=24f$ and hence $5b=f.$ By Table 2, $S$ possesses a unipotent character $\chi\neq{\mathrm{St}}_{S}$ with $\chi(1)_{p}\geq p^{91b}.$ By Lemma 3.2$(vi),$ $91b\leq 16f=80b,$ a contradiction.
@@
$(a)$ Case $S\cong{\mathrm{L}}_{n}^{\epsilon}(p^{b}),$ with $n\geq 2.$ Here $bn(n-1)=72f.$ If $n=2$ then $b=36f$ so $S={\mathrm{L}}_{2}(q^{36})$ and hence $S$ has a character of degree $q^{36}+1.$ However this is impossible as this number does not divide the order of $H.$ Hence we assume that $n\geq 3.$ If $(n,q)=(6,2)$ or $(3,2),$ then $bn(n-1)\neq 72f$. Thus $\ell_{\epsilon bn}(p)\in\pi(S)\subseteq\pi(H)$ always exists by Lemma 2.7, and so $bn\leq 18f.$ Multiplying both sides by $n-1$ and simplifying, $n\geq 5.$ By Table 2, $S$ possesses a unipotent character $\chi\neq{\mathrm{St}}_{S}$ with $\chi(1)_{p}=p^{b(n-1)(n-2)/2}.$ By Lemma 3.1(vi), $b(n-1)(n-2)/2\leq 25f.$ Multiplying both sides of the latter inequality by $n$ and simplifying, $n\leq 6.$ If $n=5,$ then $5b=18f$ and hence $4b=72f/5>14f$ so $\ell_{4p}(p)\in\pi(S)\subseteq\pi(H)$ exists and thus $4b\mid 18f=5b,$ which is impossible. Similarly if $n=6,$ then $5b=12f$ and hence $6b=72f/5>14f$ so $\ell_{6p}(p)\in\pi(S)\subseteq\pi(H)$ exists and thus $6b\mid 18f.$ Hence $12b\mid 36f=15b,$ a contradiction.
@@
$(b)$ Case $S\cong{\mathrm{S}}_{2n}(q_{1}),$ or ${\mathrm{O}}_{2n+1}(q_{1}),$ with $q_{1}=p^{b},$ $n\geq 2$ and $S\neq{\mathrm{S}}_{4}(2).$ Now $bn^{2}=36f.$ If $p=2$ and $2nb=6,$ then $n=3,b=1$ and so $bn^{2}=9=36f,$ a contradiction. Thus $\ell_{2bn}(p)\in\pi(S)\subseteq\pi(H)$ exists. Hence $2bn\leq 18f.$ Multiplying both sides by $n$ and simplifying, $n\geq 4.$ By Table 2, $S$ possesses a nontrivial irreducible character $\chi\neq{\mathrm{St}}_{S}$ with $\chi(1)_{p}\geq p^{bn(n-2)}$. By Lemma 3.1$(vi)$, $bn(n-2)\leq 25f.$ Multiplying both sides by $n$ and simplifying, $11n\leq 72$ so $n\leq 6.$ If $n=4,$ then $4b=9f$. Since $6b=54f/4>12f$, $\ell_{6p}(p)\in\pi(S)\subseteq\pi(H)$ exists and thus $6b\mid 18f=8b,$ which is impossible. Similarly if $n=5,$ then $25b=36f$. Since $10b=360f/25>14f$, $\ell_{10p}(p)\in\pi(S)\subseteq\pi(H)$ exists and thus $10b\mid 18f.$ It follows that $20b\mid 36f=25b,$ which is impossible. Thus $n=6$ and so $b=f.$ By [4], $S$ possesses a unipotent character $\chi^{\alpha}$ labeled by the symbol $\alpha=\binom{0~{}1~{}6}{~{}-~{}}$ with degree $\chi^{\alpha}(1)=(q^{6}-1)(q^{6}-q)/(2(q+1)).$ Obviously $\chi^{\alpha}(1)$ is nontrivial and different from $|H|_{p}$ and furthermore it is coprime to $\ell_{i}$ for $i=1,2,3,$ contradicting Lemma 3.1$(i)$.
@@
$(c)$ Case $S\cong{\mathrm{O}}_{2n}^{\epsilon}(q_{1}),$ where $q_{1}=p^{b},$ and $n\geq 4.$ Here $bn(n-1)=36f.$ By Table 2, $S$ possesses a unipotent character $\chi\neq{\mathrm{St}}_{S}$ with $\chi(1)_{p}\geq p^{b(n-1)(n-2)}.$ By Lemma 3.1$(vi)$, $b(n-1)(n-2)\leq 25f.$ Multiplying both sides by $n,$ and simplifying, $n\leq 6.$ If $n=4,$ then $b=3f$ and hence $q_{1}=q^{3}.$ By [4], $S$ possesses a unipotent character $\varphi$ labeled by the symbol $\binom{1~{}n}{0~{}1}$ when $\epsilon=+$ and $\binom{0~{}1~{}n}{~{}~{}1~{}~{}}$ when $\epsilon=-$ of degree
$(q_{1}^{2n}-q_{1}^{2})(q_{1}^{2}-1)=(q_{1}^{8}-q_{1}^{2})/(q_{1}^{2}-1)=q^{6}\Phi_{9}\Phi_{18}.$
It follows that $\varphi(1)\in{\mathrm{cd}}(L)-\{1,q^{36}\}$ and $(\varphi(1),\ell_{2}\ell_{3})=1.$ By Lemma 3.1$(ii)$, $\varphi(1)$ is one of the degrees in $(ii).$ Both cases are impossible by comparing the power on $q.$ If $n=5,$ then $5b=9f$ and hence $8b=72f/5>14f$ so $\ell_{8p}(p)\in\pi(S)\subseteq\pi(H)$ exists and thus $8b\mid 18f=10b,$ which is impossible. Thus $n=6$ and so $5b=6f.$ As when $n=4,$ $S$ possesses a unipotent character $\varphi$ of degree
$(q_{1}^{12}-q_{1}^{2})/(q_{1}^{2}-1)=q_{1}^{2}(q_{1}^{10}-1)/(q_{1}^{2}-1)=q_{1}^{2}(q^{12}-1)/(q_{1}^{2}-1).$
It follows that $\varphi(1)\in{\mathrm{cd}}(L)-\{1,q^{36}\}$ and $(\varphi(1),\ell_{1}\ell_{3})=1,$ contradicting Lemma 3.1$(iii)$.
@@
$(d)$ Case $S\cong{\rm G}_{2}(p^{b}),$ where $b\geq 1.$ Here $6b=36f.$ By [5], $S$ possesses an irreducible character of degree $p^{6b}-1=q^{36}-1.$ However $q^{36}-1$ does not divide $|H|,$ a contradiction.
@@
$(e)$ Case $S\in\{{}^{2}{\rm B}_{2}(q_{1}^{2}),{}^{2}{\rm G}_{2}(q_{1}^{2}),$ where $q_{1}^{2}=2^{2n+1},3^{2n+1},$ respectively with $n\geq 1.$ In these cases, the equation $|S|_{p}=|H|_{p}$ cannot occur.
@@
$(f)$ Case $S\cong{}^{2}{\rm F}_{4}(q_{1}^{2}),$ where $q_{1}^{2}=2^{2n+1},$ with $n\geq 1.$ Then $12(2n+1)=36f.$ Now $\ell_{12(2n+1)}(2)\in\pi(S)\subseteq\pi(H),$ which implies that $12(2n+1)=36f\leq 18f,$ a contradiction.
@@
$(i)$ Case $S\cong{}^{3}{\rm D}_{4}(p^{b})$. Then $12b=36f.$ Now $\ell_{12b}(p)\in\pi(S)$ but $\ell_{12b}(p)=\ell_{36f}(p)\not\in\pi(H)$, a contradiction.
@@
$(k)$ Case $S\cong{\rm E}_{7}(p^{b})$. Then $63b=36f$ and hence $7b=4f.$ By Table 1, $S$ possesses a unipotent character $\varphi\neq{\mathrm{St}}_{S}$ with $\varphi(1)_{p}=p^{46b}.$ By Lemma 3.1$(vi)$, $46b\leq 25f,$ which implies that $184b=4\cdot 46b\leq 100f=175b,$ which is impossible.
$(l)$ Case $S\cong{\rm E}_{8}(p^{b})$. Then $120b=36f$ and hence $10b=3f.$ By Table 1 and Lemma 3.1$(vi)$, $91b\leq 25f.$ It follows that $273b\leq 25\cdot 3f=250b,$ a contradiction.
$(m)$ Case $S\cong{\rm E}_{6}(p^{b})$. Then $36b=36f$ and hence $b=f.$ By [4, 13.9], $S$ has a unipotent character $\varphi$ labeled by the symbol $\phi_{6,1}$ with degree $q\Phi_{8}\Phi_{9}.$ Now $\varphi(1)\in{\mathrm{cd}}(L)-\{1,|H|_{p}\}$ is coprime to $\ell_{1}\ell_{2}\ell_{4},$ contradicting Lemma 3.1$(i)$.
The arguments for the remaining simple groups ${\rm E}_{6}(q),{\rm E}_{7}(q)$ and ${\rm E}_{8}(q)$ are similar, so we skip the details. ∎